natalie spent $135 on new shorts and tops. the shorts were 15% off and the tops were 20% off. if she saved a total of $24, how much did she spend on tops
Answer by richwmiller(17219)

s= cost of shorts
t= cost of tops
.85s+.80t=135
.15s+.20t=24
s=156 t=3
156 original total on shorts
3 original total on tops
.85*156=132.60 discounted price on shorts
.80*3=4/5*3=12/5=$2.40 discounted price on tops
.15*156+.2*3=24
23.40+.60=$24.00
She spent $ 2.40 on tops
She spent $132.60 on shorts
Another way
.85s+.80t=135
.15s+.20t=24
.85s+.80t=135
.60s+.80t=96
subtract
.25s=39
25s=3900
s=3900/25
s=156
etc